OpenStudy (anonymous):

how to solve this 13/6 - 5/18 * 3 = ?

OpenStudy (anonymous):

i got 5.6

OpenStudy (anonymous):

(13/6)-(5/18) then multiply all of this by 3

OpenStudy (anonymous):

13/6 - 5/18 * 3 - 5/18 * 3= -15/18 or -5/6 13/6 - 5/6=8/6 = 4/3

OpenStudy (anonymous):

r u shure @grazes?

OpenStudy (anonymous):

Oh. I see what you did wrong. @unmathguy , you didn't follow the order of operations

OpenStudy (anonymous):

right again!!!@grazes 4/3 is the right answer!